OpenStudy (anonymous):

Solve. Please show all of your work. A square platform has a perimeter which is 9 times the length of a side, decreased by 15. Find the length of a side.

OpenStudy (anonymous):

s = the length of a side 4s = 9s - 15 +15 = 9s - 4s 15 = 5s s = 15/5 s = 3 units Check: 4(3) = 9(3) - 15 12 = 27 - 15 did i do this correctly??

OpenStudy (paxpolaris):

yes!

OpenStudy (anonymous):

thanks!!!!!!!!!!!!!!